# Position: Scaling LLM Agents Requires Asymptotic Analysis with LLM Primitives

### v1

@@ 2 Asymptotic Analysis with LLM Primitives
Let us first define the central object of interest:
@@ 3 Example Asymptotic Separations
The three example problems in this section highlight different kinds of opportunities that would arise from having a solid complexity theory for LLM agents. The first highlights the criticality of understanding where to focus LLM size optimization in many-agent systems; the second highlights critical pitfalls of naïve anthropomorphic multi-agentization; and the third highlights issues that arise from applying LLMs to large and creative optimization tasks. Note that the analysis in this section is intentionally loose, and intended to be as rudimentary as possible, as the goal is simply to highlight the kinds of insights that can emerge, and argue that further development of AALPs will be critical in understanding how to scale real-world systems. To that end, we invite the reader to critique any assumption (implicit or explicit) made in the following analyses, with the hope that such critique will lead to more practical AALPs methodologies and thus more powerful LbAs.
@@ 3.1.2 Delegator and Specialists Approach
Suppose instead that we have a system consisting of a delegator $M_{d}$ of size $m_{d}$ and $k$ specialists $M_{1},\ldots,M_{k}$, one for each of the $k$ distinct tasks. For simplicity, assume the specialists are all of equal size $m_{s}$. Suppose the delegator can determine which of the $k$ tasks the input string belongs to by observing only a constant number of metadata tokens prepended to the input. Then, the cost of identifying the task and solving it with the designated specialist is
@@ 3.1.3 Implications
We can safely assume $m_{d}\leq m_{g}$, since solving all $k$ tasks almost certainly requires (implicitly or explicitly) identifying which task is being solved as a sub-capability. So, as $n$ grows, the comparison of interest is $\Theta(m_{g}n)$ vs. $\Theta(m_{s}n)$, i.e., $m_{g}$ vs. $m_{s}$. From analysis of the neural scaling laws of LLMs (Kaplan et al., 2020), we can assume the required size of the LLM scales with the number of required capabilities. For example, let us assume that in the present setting, the required size of the LLM scales linearly with the number of tasks it can solve. Then,
@@
Importantly, notice that this analysis shows that, for those developing the LLMs themselves, it should be much more impactful to focus on minimizing the size of the specialist LLMs than minimizing the size of the delegator. This is because, even if the delegator is relatively huge, it is called so few times relative to the specialists. In a real-world team of human developers, without such analysis, equal research and development resources might be allocated to both reducing the size of the delegator and reducing the size of specialists. For example, it could be very tempting to try to develop a relatively tiny classifier as the delegator, since it might seem like a relatively simple classification problem. Analysis like the above can preempt research on shrinking the delegator: It would be okay to use a giant off-the-self maximum-capability LLM as the delegator, and focus on shrinking the specialists. This observation supports the more general idea that, in an optimal LLM ecology, we should expect orders of magnitude more executions of smaller models than of larger models (Nisioti et al., 2024).
@@ 3.2.1 Naïve Multi-agent Approach
In this approach, we alternate between $M_{q}$ (of size $m_{q}$) identifying a bug and $M_{d}$ (of size $m_{d}$) fixing it. Suppose $M_{q}$ has the capacity to identify exactly one bug at a time, and $M_{d}$ has the capacity to fix exactly one identified bug. In the standard approach, the entire conversation history, including prior versions of the code are fed as input to each agent at each iteration. From an asymptotic perspective, such accumulation in the input may seem clearly inefficient, but this is an approach taken in popular multi-agent frameworks today, such as autogen (Wu et al., ). Suppose the QA engineer identifies bugs by producing constant size messages, while the debugging specialist rewrites the entire code every time it fixes a bug (as is also standard in current systems). Then, for the $i$th bug, the QA agent looks at all $i$ previous versions of the code and previous QA comments, yielding a total input length of $\Theta(in)$. So, the total cost of QA over all $b$ bugs is
@@ 3.2.2 Focused Multi-agent Approach
Let’s suppose all functions in the code have completely correct, precise and constant size specification (e.g., in their docstring). Now, instead of having the QA agent look at the entire code every iteration, suppose it looks only at a single function. Then, it will overall look at the entire code only a single time, resulting in a cost of
@@
Whenever a bug is identified in a function, the debugging expert then fixes the bug by looking at and rewriting only that single function, yielding a total debugging expert cost of
@@ 3.4 Exercises
  • Writing and refining a paper for submission to ICML.
@@
  • Modifying the analyses in Section 3 with the assumption that the cost of an LLM forward pass is quadratic in the input size instead of linear.
@@ 4 Alternative Views
We believe this is currently the strongest case against AALPs. The view here is that progress is being made so fast at the core of LLMs that we are bound to hit a level of general intelligence and automated AI research capabilities beyond that of human AI researchers before we need to scale modular and compute-optimized many-agent systems in the real world. Once such superintelligence is achieved, the superintelligent AI itself will be capable of performing asymptotic analysis as it sees fit, and will be otherwise responsible for developing future algorithmic insights leading to scaled agentic systems. As there is some evidence of slowing progress at the core of LLM development (Cyran, 2024; Booth, 2024), and there is always uncertainty of whether a particular research direction will hit a fundamental wall, we believe that the orthogonal AALPs approach, which foregrounds the advantages of many-agentness, is a worthwhile complement to the direct approach of LLM improvement. We also believe that since AALPs is such a fundamental tool for understanding the behavior of agentic LLM-based systems, and agentic decomposition is fundamentally essential to the optimal use of resources in the limit (i.e., a self-optimized superintelligence will not use its full power to do single-digit addition), such a superintelligence must rely on AALPs as it decides how to optimize its usage of computational resources in the physical universe, and, for those human researchers developing and applying AALPs methods, having a superintelligence use your conceptual technology is not such a bad academic legacy.
@@ 5.4 Extensions to Other Modalities
This paper argues that AALPs is required to scale LLM agents, but as agentic systems naturally grow alongside AI models that incorporate a greater and greater range of modalities (Team et al., 2023; Liang et al., 2024), the capabilities afforded by these modalities will naturally be incorporated into agentic systems; this process has already begun (Jiang et al., 2024; Gao et al., 2024; Sarch et al., 2024). The same techniques used for AALPs will naturally extend to other modalities, by encapsulating the usage of each modality as a primitive with an assigned cost. A simple implication could be that many visual tasks (when performed at scale) should be executed by compact vision-only models instead of full-fledged multi-modal foundation models.
@@ 6 Conclusion
This paper presents work whose goal is to advance the field of artificial intelligence. There are many societal consequences that potentially follow from such work. One main advantage of the approach advocated in this paper is the potential reduction in future machine suffering, as discussed in Section 5.6. Another advantage of the fact that AALPs leads naturally to maximally decomposed problem-solving is that decomposed systems can be much more interpretable, auditable, and therefore safer than a single opaque centralized AI carrying out the same task, a point that has been acknowledged in prior work (Khot et al., 2021; Sharkey et al., 2025).

### v2

@@ 2 Asymptotic Analysis with LLM Primitives
The premise of asymptotic analysis is that by focusing on counting the executions of the fundamental operations whose cost we most care about, we can characterize the fundamental trends in cost as the system scales. Such analysis can reveal scaling insights that might be obscured by implementation details or cost overheads that dominate at the smaller scales that are empirically viable today.
Let us first define our central object of interest:
@@
Then, if LLMs themselves are developed in a way that alters their cost and/or capabilities, these results can be directly incorporated into AALPs to identify implications for particular LbAs. AALPs also provides a framework for analyzing LbAs that require capabilities that LLMs are not yet capable of, thereby allowing researchers to be more prepared for the future when it arrives.
@@ 3 Example Asymptotic Separations
$k$-Task (Sec. 3.1) Debugging (Sec. 3.2) Evolution (Sec. 3.3)
Agents Generalist $M_{g}$ Delegator $M_{d}$ Specialists $M_{s}^{1},\ldots M_{s}^{k}$ QA Engineer $M_{q}$ Software Developer $M_{d}$ Mutator $M_{u}$
Optimistic $y\leftarrow M_{g}(x)$ for $i\in 0\ldots b-1$ do     $\textrm{bug}\leftarrow M_{q}(x)$     $x\leftarrow M_{d}(\textrm{bug},x)$ while $f(x)<k$ do     $x^{\prime}\leftarrow M_{u}(x)$     if $f(x^{\prime})>f(x)$ then          $x\leftarrow x^{\prime}$
Optimized $i\leftarrow M_{d}(x_{0:c})$ $y\leftarrow M_{s}^{i}(x)$ for $i\in 0\ldots k-1$ do     $\textrm{bug}\leftarrow M_{q}(x_{il:il+l})$     if bug then          $x_{il:il+l}\leftarrow M_{d}(\textrm{bug},x_{il:il+l})$ $i\leftarrow 0$ while $f(x)<k$ do     $x^{\prime}\leftarrow x$     $x^{\prime}_{il:il+l}\leftarrow M_{u}(x_{il:il+l})$     if $f(x^{\prime})>f(x)$ then          $x\leftarrow x^{\prime}$          $i\leftarrow i+1$
Improvement $\Theta(k)$ $\Theta(bk^{2})$ $\Theta(2^{k}k)$
Table 1: Overview of Examples. This table gives a high-level summary of the examples described in Section 3. It lists the LLM agents used, and gives pseudocode for the optimistic implementation (i.e., based on an intuitive belief in the power of LLMs) and the optimized one (based on a more careful LbA design (Def. 2.1)). The improvements result from applying AALPs to these implementations.
@@
The three example problems in this section highlight different kinds of opportunities that would arise from having a solid complexity theory for LLM agents. The first highlights the criticality of understanding where to focus LLM size optimization in many-agent systems; the second highlights critical pitfalls of naïve anthropomorphic multi-agentization; and the third highlights issues that arise from applying LLMs to large and creative optimization tasks. See Table 1 for a high-level summary. Note that the analysis in this section is intentionally loose, and intended to be as rudimentary as possible, as the goal is simply to highlight the kinds of insights that can emerge, and argue that further development of AALPs will be critical in understanding how to scale real-world systems. To that end, we invite the reader to critique any assumption (implicit or explicit) made in the following analyses, with the hope that such critique will lead to more practical AALPs methodologies and thus more powerful LbAs.
@@ 3.1.2 Delegator and Specialists Approach
Suppose instead that we have a system consisting of a delegator $M_{d}$ of size $m_{d}$ and $k$ specialists $M_{s}^{1},\ldots,M_{s}^{k}$, one for each of the $k$ distinct tasks. Several recent approaches take this kind of LLM-routing approach (Hu et al., 2024; Feng et al., 2025; Narayan et al., 2025). For simplicity, assume the specialists are all of equal size $m_{s}$. Suppose the delegator can determine which of the $k$ tasks the input string belongs to by observing only a constant number of metadata tokens prepended to the input, e.g., the delegator has been fine-tuned so that descriptions of the $k$ tasks need not be supplied in the prompt. Then, the cost of identifying the task and solving it with the designated specialist is
@@ 3.1.3 Implications
We can safely assume $m_{d}\leq m_{g}$, since solving all $k$ tasks almost certainly requires (implicitly or explicitly) identifying which task is being solved as a sub-capability. [^1] Evidence for this can be found in mixture-of-experts models: In Mixtral 8x7B (Jiang et al., 2024a), each expert (analogous to a specialist) has $\approx$7B parameters, while the router (delegator) only has $\approx$32K, and the overall system outperforms a generalist LLM with $\approx$70B parameters; DeepSeek-V3 (Liu et al., 2024) routes across 256 experts, by simply comparing a single input token embedding to the “centroid” token embedding of each expert. So, as $n$ grows, the comparison of interest is $\Theta(m_{g}n)$ vs. $\Theta(m_{s}n)$, i.e., $m_{g}$ vs. $m_{s}$. From analysis of the neural scaling laws of LLMs (Kaplan et al., 2020), we can assume the required size of the LLM scales with the number of required capabilities. For example, let us assume that in the present setting, the required size of the LLM scales linearly with the number of tasks it can solve. Then,
@@
Importantly, notice that this analysis shows that, for those developing the LLMs themselves, it should be much more impactful to focus on minimizing the size of the specialist LLMs than minimizing the size of the delegator. This is because, even if the delegator is relatively huge, it is called so few times relative to the specialists, and the cost of any initial fine-tuning of the delegator is amortized over the lifetime of the system. In a real-world team of human developers, without such analysis, equal research and development resources might be allocated to both reducing the size of the delegator and reducing the size of specialists. For example, it could be very tempting to try to develop a relatively tiny classifier as the delegator, since it might seem like a relatively simple classification problem. Analysis like the above can preempt research on shrinking the delegator: It would be okay to take a giant off-the-self maximum-capability LLM as the delegator, fine-tune it once, and then focus on shrinking the specialists. This observation supports the more general idea that, in an optimal LLM ecology, we should expect orders of magnitude more executions of smaller models than of larger models (Nisioti et al., 2024).
@@ 3.2.1 Naïve Multi-agent Approach
In this approach, we alternate between $M_{q}$ (of size $m_{q}$) identifying a bug and $M_{d}$ (of size $m_{d}$) fixing it. Suppose $M_{q}$ has the capacity to identify exactly one bug at a time, and $M_{d}$ has the capacity to fix exactly one identified bug. In the standard approach, the entire conversation history, including prior versions of the code are fed as input to each agent at each iteration. From an asymptotic perspective, such accumulation in the input may seem clearly inefficient, but this is an approach taken in popular multi-agent frameworks today, such as autogen (Wu et al., ). Suppose the QA engineer identifies bugs by producing constant size messages, while the debugging specialist rewrites the entire code every time it fixes a bug (as is also standard in current systems). Then, for the $i$th bug, the QA agent looks at all $i$ previous versions of the code and previous QA comments, yielding a total input length of $\Theta(in)$. So, the cost of QA over all $b$ bugs is
@@ 3.2.2 Focused Multi-agent Approach
Let’s suppose all functions in the code have completely correct, precise and constant size specification (e.g., in their docstring). This implies the full code can be fixed by fixing each function independently. Now, instead of having the QA agent look at the entire code every iteration, suppose it looks only at a single function. Then, it will overall look at the entire code only a single time, resulting in a cost of
@@
Whenever a bug is identified in a function, the debugging expert then fixes the bug by looking at and rewriting only that single function, yielding a total bug-fixing cost of
@@ 3.4 Exercises
  • Writing and refining a paper for submission to a conference dedicated to AI-produced AI research.
@@ 4 Alternative Views
We don’t need LLM primitives: Existing cost measures are enough—such as simply counting tokens, GPU hours, or floating point operations.
This view raises the question of what level of granularity is most useful for understanding LbA behavior. If we count tokens without regard to model size, we miss out on the increasing differences in model scale required for different capabilities; if we count GPU hours, total activated LLM parameters, or floating point operations, the results will be conditional on the particular hardware and LLM internal implementations available at the time. By encapsulating model cost and capabilities, AALPs provides the right level of abstraction for reasoning about the costs of scaled agentic systems, orthogonal to the ongoing development of LLMs.
Alternative View 4:
@@
Prominent AI researchers are preparing for the possibility of super-intelligence within a few years (Grace et al., 2024; Altman, 2024; Hendrycks et al., 2025; Fortson, 2025). We believe this presents the strongest case against AALPs. The view here is that progress is being made so fast at the core of LLMs that we are bound to hit a level of general intelligence and automated AI research capabilities beyond that of human AI researchers before we need to scale modular and compute-optimized many-agent systems in the real world. Once such superintelligence is achieved, the superintelligent AI itself will be capable of performing asymptotic analysis as it sees fit, and will be otherwise responsible for developing future algorithmic insights leading to scaled agentic systems. As there is some evidence of slowing progress at the core of LLM development (Cyran, 2024; Booth, 2024), and there is always uncertainty of whether a particular research direction will hit a fundamental wall, we believe that the orthogonal AALPs approach, which foregrounds the advantages of many-agentness, is a worthwhile complement to the direct approach of LLM improvement. We also believe that since AALPs is such a fundamental tool for understanding the behavior of agentic LLM-based systems, and agentic decomposition is fundamentally essential to the optimal use of resources in the limit (i.e., a self-optimized superintelligence will not use its full power to do single-digit addition), such a superintelligence must rely on AALPs as it decides how to optimize its usage of computational resources in the physical universe, and, for those human researchers developing and applying AALPs methods, having a superintelligence use your conceptual technology is not such a bad academic legacy.
@@ 5 Research Directions
5.1 Identifying Useful Assumptions
As alluded to throughout this paper, AALPs provides a template for systematically reasoning about the efficiency of LLM-based agentic systems, but any specific application of AALPs depends on the particular assumptions adopted, and which assumptions are most useful for driving progress is an open question. For example, are the linear scaling assumptions in Definition 2.2 and Section 3.1.3 reasonable? Is it reasonable to disregard relatively large constants like the system prompt and explanations surrounding an answer, or assume they can be distilled away? Is it reasonable to assume functional independence of solution components as is done in Sections 3.2 and 3.3, or should we explicitly deal with approximations in cases of messy problem decompositions or in the absence of an oracle? Moving forward, it will be important for AALPs to handle more nuanced assumptions based on complexities arising in practice.
@@ 5.5 Extensions to Other Modalities
This paper argues that AALPs is required to scale LLM agents, but as agentic systems naturally grow alongside AI models that incorporate a greater and greater range of modalities (Team et al., 2023; Liang et al., 2024), the capabilities afforded by these modalities will naturally be incorporated into agentic systems; this process has already begun (Jiang et al., 2024b; Gao et al., 2024; Sarch et al., 2024). AALPs does not immediately extend to other modalities, due to its basis in LLM primitives, but similar techniques could be used to encapsulate the usage of each modality as a primitive with an assigned cost. A simple implication could be that many visual tasks (when performed at scale) should be executed by compact vision-only models instead of full-fledged multi-modal foundation models.
@@ 6 Conclusion
Acknowledgments
We would like to thank the Cognizant AI Lab research team for useful discussions in the development of this work, and, in particular, Dan Fink and Giuseppe Paolo for early inspiring discussions before this work started to take form.
@@ Acknowledgments
This paper presents work whose goal is to advance the field of artificial intelligence. There are many societal consequences that potentially follow from such work. One main advantage of the approach advocated in this paper is the potential reduction in future machine suffering, as discussed in Section 5.7. Another advantage, arising from the fact that AALPs leads naturally to maximally decomposed problem-solving, is that decomposed systems can be much more interpretable, auditable, and therefore safer than a single opaque centralized AI carrying out the same task, a point that has been acknowledged in prior work (Khot et al., 2021; Sharkey et al., 2025).